# Accelerating Molecular Dynamics Simulations with Foundation Neural Network Models using Multiple Time-Step and Distillation

### v1

@@ Abstract
We present a strategy to accelerate molecular dynamics simulations using foundation neural network models. To do so, we apply a dual-level neural network multi-time-step (MTS) strategy where the target accurate potential is coupled to a simpler but faster model obtained via a distillation process. Thus, the 3.5 Å-cutoff distilled model is sufficient to capture the fast varying forces, i.e. mainly bonded interactions, from the accurate potential allowing its use in a reversible reference system propagator algorithms (RESPA)-like formalism. The approach conserves accuracy, preserving both static and dynamical properties, while enabling to evaluate the costly model only every 3 to 6 fs depending on the system. Consequently, large simulation speedups over standard 1 fs integration are observed: 4-fold in homogeneous systems and 2.3-fold in large solvated proteins. Such a strategy is applicable to any neural network potential and reduces their performance gap with classical force fields.
@@ 3.2 Solvated Molecules
To further assess the accuracy of the proposed MTS scheme, we turn to small solvated molecules. We observe stability limits similar to those for bulk water for the five molecules that we tested: ethanol, benzene, trimethylamine, diethylsulfide, and acetic acid. We then evaluated the hydration free energies (HFEs) of these (as well as water). Calculations were performed leveraging the alchemical Lambda-ABF 45 method, employing both system-specific models trained on-the-fly and a generic small foundation model. All simulations were carried out using the same MTS integration scheme described above, with an inner time step of and an outer time step of .
The predicted HFEs were compared to the STS reference values, as shown in Figure 3. The system-specific model achieved a mean absolute error (MAE) of , a root-mean-square error (RMSE) of , and a coefficient of determination () of . The generic foundation model also showed good performance, with an MAE of , an RMSE of , and of .
@@ 3.3 Protein-ligand complexes
Using an inner time step of and an outer time step of with HMR, stable simulations over were obtained while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5. This setup provides a speedup of 2.3 with respect to the STS simulation with time step, strongly accelerating production simulations with biological systems.
@@ 4 Conclusion
We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.3 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties. Note that this is a first estimation of the computational gains as the code has yet to be optimized further to handle the dual-level approach more efficiently. All in all, our approach allows to reach nearly of simulation per day on a single A100 GPU for a realistic protein-ligand complex while preserving the ab initio-like accuracy of the FeNNix-Bio1(M) model. In combination with well-established accelerated sampling schemes 47, 48, 49, this enables truly large-scale simulations with foundation neural neural network potentials.

### v2

@@ Abstract
We present a strategy to accelerate molecular dynamics simulations using foundation neural network models. To do so, we apply a dual-level neural network multi-time-step (MTS) strategy where the target accurate potential is coupled to a simpler but faster model obtained via a distillation process. Thus, the 3.5 Å-cutoff distilled model is sufficient to capture the fast varying forces, i.e. mainly bonded interactions, from the accurate potential allowing its use in a reversible reference system propagator algorithms (RESPA)-like formalism. The approach conserves accuracy, preserving both static and dynamical properties, while enabling to evaluate the costly model only every 3 to 6 fs depending on the system. Consequently, large simulation speedups over standard 1 fs integration are observed: 4-fold in homogeneous systems and 2.7-fold in large solvated proteins. Such a strategy is applicable to any neural network potential and reduces their performance gap with classical force fields.
@@ 3.2 Solvated Molecules
To further assess the accuracy of the proposed MTS scheme, we turn to small solvated molecules. We observe stability limits similar to those for bulk water for the five molecules that we tested: ethanol, benzene, trimethylamine, diethylsulfide, and acetic acid. We then evaluated the hydration free energies (HFEs) of these (as well as water). Calculations were performed leveraging the alchemical Lambda-ABF 45 method, employing both system-specific models trained on-the-fly and a generic small foundation model. All simulations were carried out using the same MTS integration scheme described above, with an inner time step of and an outer time step of and HMR. For benzene MTS simulations using the generic model, we needed to reduce the outer time step to to avoid instabilities. This confirms that the system specific model is more robust than the generic one.
Figure 3 shows the predicted HFEs compared to the STS reference values. The system-specific model achieved a mean absolute error (MAE) of , a root-mean-square error (RMSE) of , and a coefficient of determination () of . The generic foundation model also showed good performance, with an MAE of , an RMSE of , and of .
@@ 3.3 Protein-ligand complexes
Using an inner time step of and an outer time step of with HMR, we obtain stable simulations over while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5. This setup provides a speedup of 2.7 with respect to the STS simulation with time step, strongly accelerating production simulations with biological systems.
@@ 4 Conclusion
We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.7 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties. Note that this is a first estimation of the computational gains as the code has yet to be optimized further to handle the dual-level approach more efficiently. All in all, our approach allows to reach nearly of simulation per day on a single A100 GPU for a realistic protein-ligand complex while preserving the ab initio-like accuracy of the FeNNix-Bio1(M) model. In combination with well-established accelerated sampling schemes 47, 48, 49, this enables truly large-scale simulations with foundation neural neural network potentials.